# Comparative Analysis of Weather-Based Indexes and the Actuaries Climate IndexTM for Crop Yield Prediction

### v1

@@ 3.2 Second Stage
The predictive capabilities of climate indexes on the yields of corn (grain), wheat (winter), and soybeans are examined using both traditional statistical approaches and modern machine learning algorithms. In the literature, regression models (Shastry et al. 2017; Shah et al. 2018; Ansarifar et al. 2021) and machine learning techniques (Van Klompenburg et al. 2020; Nigam et al. 2019; Iniyan et al. 2023; Mahesh and Soundrapandiyan 2024) are frequently used for crop yield prediction, primarily due to their simplicity, interpretability, and computational efficiency. In this study, we use GLM and GAM, which are favored for their flexibility regarding distributional assumptions (detailed in Appendix A.1). Furthermore, to demonstrate the advantages of machine learning algorithms over traditional methods, we incorporate two widely used gradient boosting algorithms, XGB and LGBM (detailed in Appendix A.2).
@@
Crop yields are influenced not only by weather conditions during the harvest period but also by extreme weather events throughout the growing season. Such conditions, including temperature extremes and precipitation anomalies, can have significant cumulative effects on crop development and productivity. To capture the full spectrum of potential impacts, all seasonal values for each index are included in the data sets. Additionally, the harvest period of each crop is carefully considered when structuring the data sets. Since corn and soybeans are typically harvested in autumn months, while wheat is harvested in summer months, the seasonal climate indexes are aligned accordingly. For corn and soybeans, the data sets cover the period from winter to autumn, whereas for wheat, they span from autumn to summer for each year. This approach ensures that the data sets encompass the entire year leading up to each crop’s harvest period. The analyses fit a total of 22 models, each incorporating explanatory variables representing various weather measurements and their combinations. A detailed breakdown of each model and the associated data sets is presented in Table 2. Comprehensive information on the data sets, along with the crop yields fitted by each model, is provided in full in Section 4.
@@ 22 ACI
Note. Based on four meteorological season values for each index.
The choice of explanatory variables for each model follows a logical progression, designed to evaluate the predictive power of various index configurations on crop yields. The first seven models are composed of WBI components and their combinations. The next set of seven models (Models 8–14) features indexes based on ACI components that correspond on a one-to-one basis with WBI variables. Additional ACI components (W, D, S) that are not represented in the initial 14 models are incorporated into the third set of seven models (Models 15–21). Lastly, the final one (Model 22) uses the composite ACI, which aggregates all six components into a single index, offering a comprehensive view of climate conditions in a composite measure.
A flowchart of the proposed methodology is illustrated in Figure 1. The methodology includes two distinct comparison frameworks to evaluate model performances based on different input variables. In the first comparison analysis, the performance of regression models and machine learning algorithms is assessed using principal components derived from the explanatory variables. Since all indexes incorporated into each model are weather-based, they tend to reflect similar meteorological patterns, which may result in multicollinearity issues in regression analyses. To ensure independence between the explanatory variables, and simplify the explanatory data sets without compromising model accuracy, we use PCA and FPCA (see details in Appendix A.3) as dimensionality reduction techniques. By comparing these two techniques, this study also evaluates the performance of FPCA relative to PCA in modeling time-dependent data like climate indexes.
Unlike regression models, machine learning algorithms do not require data simplification or the elimination of multicollinearity among variables. Therefore, in the second comparison analysis, the climate indexes are directly used as input variables in the machine learning algorithms without applying any dimensionality reduction. This approach not only allows for a direct assessment of model performance using the raw climate data but also provides a basis for evaluating the effectiveness of the principal component approach proposed in the methodology.
@@
The “scikit-learn” (Fabian 2011) and “scikit-fda” (Ramos-Carreño et al. 2024) Python libraries are used to implement PCA and FPCA with a B-spline basis, respectively. To maintain consistency and simplicity in model execution, the default settings of both packages are applied. From the transformed data sets, the first four principal components ($PC_{1}$, $PC_{2}$, $PC_{3}$, $PC_{4}$), which collectively account for more than 85% of the variance in the data, are selected as the input variables for analyses. This selection ensures that the models capture the majority of the information while reducing dimensionality.
@@
Before conducting the regression analyses, the next critical step involves determining the appropriate link function of GLMs and GAMs, as well as the distribution of crop yield data. The Python package “distfit” (Taskesen 2020) is used to identify the most suitable distribution. This package evaluates 89 univariate distributions and selects the best-fitting one by comparing their residual sum of squares. To ensure consistency across GLMs and GAMs, the log-link function is applied in all models. This function is particularly well-suited for crop yield data, as these values are inherently positive and continuous. The implementation of GLMs and GAMs is carried out using the R packages "stats" (Team 2022) and "mgcv" (Wood 2015), respectively. The GLM used in the analyses is expressed as
$log(E(Y|PC_{1},PC_{2},PC_{3},PC_{4}))=\beta_{0}+\sum_{i=1}^{4}\beta_{i}*PC_{i},$ (9)
whereas the GAM is as follows
$log(E(Y|PC_{1},PC_{2},PC_{3},PC_{4}))=\beta_{0}+\sum_{i=1}^{4}f_{i}(PC_{i}).$ (10)
@@
Here, $Y$ represents the crop yield data, $PC_{i}$ $(i=1,2,3,4)$ denote the principal components obtained through FPCA. $\beta_{i}$’s are the parameters, and $f_{i}(.)$ are smooth functions capturing the potentially nonlinear relationships between principle components and the crop yields. In the analyses, $Y$ and $PC_{i}$ are included as vectors of size $n\times 1$ where $n$ represents the number of observations (years).
[…]
@@ 4.1 Data
Each component of the ACI consists of a time series, either on a monthly basis, or for meteorological seasons (three months ending February, May, August, November). These time series, starting in 1961, are derived from measurements collected at meteorological stations across the U.S. and Canada. The data are constructed separately for each country, in combination, and across 12 sub-regions. For consistency, this study uses U.S. data, where states are grouped into seven sub-regions based on their geographic locations: Alaska (ALA), Central East Atlantic (CEA), Central West Pacific (CWP), Midwest (MID), Southeast Atlantic (SEA), Southern Plains (SPL), and Southwest Pacific (SWP). The analyses use standardized seasonal values for each of the six ACI components (T90, T10, P, D, W, S) as well as the composite index (ACI). These seasonal data, obtained from the official Actuaries Climate IndexTM website (https://actuariesclimateindex.org/data/, last accessed on April 18, 2025), cover a 63-year period from 1961 to 2023 for all seven sub-regions.
@@
Regarding weather-based indexes, the National Oceanic and Atmospheric Administration (https://www.ncei.noaa.gov/access/monitoring/products/, last accessed on April 18, 2025) provides monthly data on CDD, HDD, and PRE for all U.S. states, covering the period from 1895 to the present. To align with the ACI data, seasonal values for CDD, HDD, and PRE between 1961 and 2023 are calculated by summing the monthly values for each meteorological season. Sub-regional data is subsequently generated by averaging the seasonal values of all states within each region. However, due to the unavailability of CDD and HDD data for Alaska, the analyses are limited to the six remaining regions of the U.S.

### v2

@@ 3.1 Crop Yield Prediction
The predictive capabilities of climate indexes on crop yields are evaluated using both traditional statistical models and modern machine learning algorithms. In the literature, regression models (Shastry et al. 2017; Shah et al. 2018; Ansarifar et al. 2021) and machine learning techniques (Van Klompenburg et al. 2020; Nigam et al. 2019; Iniyan et al. 2023; Mahesh and Soundrapandiyan 2024; You et al. 2017; Crane-Droesch 2018) are widely used for crop yield prediction, largely due to their ease of implementation, interpretability, and computational efficiency. In this study, the GLM and GAM, which are favored for their flexibility in handling distributional assumptions (detailed in Appendix A.1), are used. To complement these approaches and highlight the advantages of machine learning methods over traditional statistical models, two widely used gradient boosting algorithms, XGB and LGBM (detailed in Appendix A.2), are also incorporated.
@@
Crop yields are influenced not only by weather conditions during the harvest period but also by extreme weather events occurring throughout the entire growing cycle, from planting to harvest. Adverse conditions such as temperature extremes and precipitation anomalies can have significant cumulative effects on crop development and productivity. To capture the full spectrum of potential impacts, the meteorological seasonal values of each index covering the entire year are included in the data sets. Additionally, the timing of harvest for each crop is carefully considered to ensure that the data sets encompass the complete year leading up to harvest, thereby reflecting all relevant weather impacts. The analyses fit a total of 22 models, each incorporating explanatory variables representing various weather measurements and their combinations. A detailed breakdown of each model and the associated data sets is presented in Table 1. Comprehensive information on the data sets, along with the crop yields fitted by each model, is provided in full in Section 4.
@@ 22 ACI
Despite differences in their calculation methods, the WBIs and ACI components can be systematically grouped based on the weather conditions they represent. Specifically, CDD (Eq. 2) and T90 (Eq. 5) both capture warm weather extremes, whereas HDD (Eq. 3) and T10 (Eq. 6) measure cold weather conditions. Similarly, PRE (Eq. 4) and P (Eq. 7) reflect the severity of precipitation. This classification provides a clear, logical, and consistent framework for comparing the indexes, ensuring a robust basis for evaluating their predictive power with respect to crop yields.
The first seven models summarized in Table 1 are built using only WBIs and their combinations, providing a baseline for comparison. Models 8-14 replace these with ACI components that correspond one-to-one with the WBIs, enabling direct performance comparisons between the two index types. To extend the analysis, additional ACI components (W, D, S) that are not represented in the first 14 models are sequentially incorporated into Models 15-21. Finally, Model 22 uses the composite ACI, which aggregates all six components into a single index, offering a comprehensive measure of climate conditions.
A flowchart of the proposed methodology is presented in Figure 1. The framework consists of two distinct comparison structures designed to evaluate model performance under different input variables. In the first comparison analysis, the performance of regression models and machine learning algorithms is assessed using principal components derived from the explanatory variables. Since all indexes incorporated into the models are weather-based, they tend to reflect similar meteorological patterns, which may result in multicollinearity issues in regression analyses. To address this and simplify the explanatory data sets without compromising model accuracy, the PCA and FPCA (see details in Appendix A.3) are used as dimensionality reduction techniques. Comparing these two techniques also allows the performance evaluation of FPCA over PCA, particularly in modeling time-dependent data such as climate indexes.
Unlike regression models, machine learning algorithms do not require prior data simplification or the elimination of multicollinearity among variables. Therefore, the second comparison analysis directly uses the raw climate indexes as input variables into the machine learning algorithms, without applying any dimensionality reduction techniques. This approach not only allows for a direct assessment of model performance using the raw climate data but also provides a basis for evaluating the effectiveness of the principal component approach proposed in the methodology.
@@
The Python packages “scikit-learn” (Fabian 2011) and “scikit-fda” (Ramos-Carreño et al. 2024) are used to implement PCA and FPCA with a B-spline basis, respectively. To maintain consistency and simplicity in model execution, the default settings of both packages are applied. From the transformed data sets, the first four principal components ($PC_{1}$, $PC_{2}$, $PC_{3}$, $PC_{4}$), which collectively account the majority of the variance in the data, are selected as the input variables for analyses.
Time-Trend Analysis
In crop yield analysis, detrending time series is essential, as agricultural production often follows a long-term upward trend driven by technological innovations, and advances in farming practices. Without appropriately removing or controlling for the trend in time, predictive models especially regression-based approaches may overstate their performance by simply capturing the persistent growth pattern instead of isolating the actual effects of weather and climate factors on yields.
A wide range of detrending techniques is available in the literature, including both parametric and non-parametric methods. Common approaches include linear trend (Zhou et al. 2021), quadratic trend (Bucheli et al. 2022), moving average smoothing (Lu et al. 2017), locally weighted regression (Lu et al. 2017), and two-knot linear spline (Liu and Ker 2021). In this study, the linear detrending method, the most widely used approach in agricultural economics, is adopted, as it offers both simplicity and interpretability. The detrended crop yields ($Dy_{t}$) are defined as:
$Dy_{t}=y_{t}-\hat{y}_{t}+\hat{y}_{0},$ (9)
where $y_{t}$ represents the observed crop yields, $\hat{y}_{t}$ denotes the values predicted from the fitted linear regression model, in which time $t$ is used as the independent variable. The addition of $\hat{y}_{0}$ ensures that all detrended yields are shifted to the level of the first predicted yield (Jewson and Brix 2005). This adjustment prevents the occurrence of negative values in the detrended series, which is particularly important because GLM and GAM with some specific link functions cannot accommodate negative-valued inputs. From Equation (9), the removed time trends ($TT_{t}$) from the observed crop yields are obtained as follows:
$TT_{t}=y_{t}-Dy_{t}=\hat{y}_{t}-\hat{y}_{0}.$ (10)
The time stationarity of crop yields is evaluated using the Augmented Dickey-Fuller (ADF) test (Dickey and Fuller 1979), where the null hypothesis indicates that the crop yield series is non-stationary.
Modeling
@@
After detrending the crop yield data, the next crucial step is to determine both the appropriate distribution of the yields and the most suitable link function for GLMs and GAMs before performing regression analyses. The Python package “distfit” (Taskesen 2020) is used to identify the best-fitting distribution. This package evaluates 89 univariate distributions and selects the optimal one based on the residual sum of squares. For the link function, consistency across GLMs and GAMs is ensured by applying the log-link function in all models. The log-link function is particularly well-suited for crop yield data, as these values are inherently positive and continuous. The implementation of GLMs and GAMs is carried out using the R packages "stats" (Team 2022) and "mgcv" (Wood 2015), respectively. The GLM and GAM are expressed as:
$log(E(Dy|PC_{1},PC_{2},PC_{3},PC_{4}))=\beta_{0}+\sum_{h=1}^{4}\beta_{h}*PC_{h},$ (11)
$log(E(Dy|PC_{1},PC_{2},PC_{3},PC_{4}))=\beta_{0}+\sum_{h=1}^{4}f_{h}(PC_{h}),$ (12)
@@
respectively. Here, $Dy$ represents the detrended crop yields, and $PC_{h}$ $(h=1,2,3,4)$ denote the principal components. The coefficients $\beta_{h}$ are the model parameters, and $f_{h}(.)$ are smooth functions that capture potentially nonlinear relationships between principal components and the crop yields. In the analyses, both $Dy$ and $PC_{h}$ are treated as column vectors of size $N\times 1$, where $N$ corresponds to the number of observations (years).
To construct the XGB and LGBM models, the Python packages “xgboost” (Brownlee 2019) and “lightgbm” (Microsoft 2017) are chosen among the alternatives, respectively. To maintain consistency, the same hyper-parameters are used in all models. These are the number of estimators (500), maximum depth (2), and learning rate (0.01).
Cross-Validation
Given the time series nature of climate indexes, training and test sets cannot be selected randomly when applying cross-validation methods. A random split would create the unrealistic situation in which future data points are used to predict past values. Therefore, we apply the iterative $M$-split leave-$k$-out cross-validation method, which is specifically suitable for time series applications. In this approach, the data set is divided into $M$ sequential splits, where each test set contains a unique block of $k$ consecutive observations. By iteratively leaving out these $k$ observations for testing, while the preceding observations are used for training, the method ensures that the natural time order is preserved. The cross-validation procedure applied in this study is illustrated in Figure 2.
Figure 2: $M$-split leave-$k$-out cross-validation method
In the procedure, $M$ is the number of splits, $N$ denotes the total number of observations (years), while $k$ is set to approximately 10% of $N$. Accordingly, the train/test share for the first iteration (split) is arranged as 90% training data and 10% testing data. Model performance is assessed on both training and test sets using the Mean Absolute Percentage Error (MAPE). The models are fitted to the detrended crop yields; however, the removed time trends are reintroduced during performance evaluation to ensure that predictive accuracy reflects the true scale of observed yields. The overall MAPE of the model across all splits for the training and test sets are:
$MAPE_{Train}=\frac{1}{M}\sum_{m=1}^{M}\bigg[\frac{1}{N-mk}\sum_{t=1}^{N-mk}\frac{|y_{t}-\hat{Dy}_{t}+TT_{t}|}{y_{t}}\bigg],$ (13)
$MAPE_{Test}=\frac{1}{M}\sum_{m=1}^{M}\bigg[\frac{1}{k}\sum_{t=N-mk+1}^{N-mk+k}\frac{|y_{t}-\hat{Dy}_{t}+TT_{t}|}{y_{t}}\bigg],$ (14)
respectively. Here, $y_{t}$ denotes the observed crop yields, $\hat{Dy}_{t}$ represents the predicted detrended crop yields, and $TT_{t}$ corresponds to the removed time trends, where $t$ indicates time (year).
3.2 Weather-Derivative Pricing
For the financial market application, the similarities between WBIs and the components of the ACI are examined by evaluating the strength of their correlations and comparing their empirical cumulative distribution function (ECDF) values. The comparison is structured across three groups, each representing a different type of weather condition, thereby enabling a systematic and detailed comparison of the indexes. Mainly, the focus is on testing the performances of indexes within the weather-derivative framework and comparing their outcomes. For this purpose, the payoffs of weather derivative contracts are analyzed using underlying variables derived from both the WBIs and ACI components. This approach permits to investigate whether the indexes exhibit similar financial performance in market-based applications, thereby extending their potential utility beyond the agricultural sector.
Index Comparison
The grouping of indexes, established to enable meaningful comparisons, is summarized in Table 2. The time periods assigned to each group (Group 1: Summer, Group 2: Winter, Group 3: Spring) are carefully selected to align with the practical applications of climate indexes. For example, CDD is widely used to estimate energy demand during the summer cooling season, whereas HDD is frequently used to assess heating requirements in the winter (Spinoni et al. 2018). Regarding precipitation, its timing plays a critical role in agricultural productivity. Wang et al. (2016) highlight that crop yields are especially sensitive to rainfall during the rainy season, whereas Peltonen-Sainio et al. (2011) show that precipitation before harvest often negatively impacts crop yields. Therefore, for the precipitation group, index comparisons are conducted for the spring months to encompass the rainy season preceding the general harvest period of most crops.
Table 2: Structure of groups
Group
Weather-Based
Indexes (WBIs)
Components
of the ACI
Meteorological
Season
Measurement
1 CDD T90 Summer Warm weather
2 HDD T10 Winter Cool weather
3 PRE P Spring Precipitation
The dependence between the indexes is analyzed using the Spearman rank correlation coefficient (Zar 2005), while distributional comparisons are conducted through the Anderson-Darling (A-D) test (Anderson and Darling 1954). The A-D test is chosen over the more commonly known Kolmogorov-Smirnov test because the former is more sensitive to the tails of the distribution (Engmann and Cousineau 2011), making it particularly suitable in applications involving climate indexes, where extreme weather conditions often play a critical role in determining outcomes.
Derivative Contracts
Weather derivatives are structured in several contract forms based on different underlying weather indexes. In the literature, the most commonly used indexes are CDD, HDD, and PRE (Solanilla Blanco 2025; Odening et al. 2007). Regarding contract types, the most frequent forms are call/put options, and swaps (Jewson and Brix 2005; Alaton et al. 2002). Basically, a call option provides a payout when the underlying weather index exceeds a predefined threshold, whereas a put option triggers a payout if the index falls below that threshold. Swaps, on the other hand, are structured as agreements between two parties to exchange cash flows based on the difference between the actual observed weather outcomes and an agreed-upon strike level. In this study, an uncapped call option is used, whose payoff $PO$ is defined as follows:
$PO=\alpha\,max\{I-K,0\},$ (15)
where $I$ represents the value of the underlying weather index, and $K$ denotes the predetermined strike level. Here, $\alpha$ is the tick size, specifying the monetary value assigned per unit of the index.
Pricing Methods
Pricing of weather derivatives involves determining a fair contract value using one of several methods. Two commonly used ones, which are also adopted in this study, are the Historical Burn Analysis (HBA) and Index Modeling (IM) (Alexandridis and Zapranis 2012). The HBA is widely used as a benchmark because of its simplicity, since it does not require fitting a probability distribution to the weather index or solving complex stochastic equations. The idea behind HBA is retrospective: it evaluates how the contract would have been performed in previous years and estimates the expected future payoff as the average of historical payoffs. The fair price of a call option under HBA can be computed with the following steps:
  1. (i)
    Determine the fair strike level ($K$) using historical weather index values.
  2. (ii)
    Calculate the historical payoffs of the option for each year based on the observed index values.
  3. (iii)
    Compute the mean of the historical payoffs, which serves as a fair price.
Alternatively, within the IM method, the underlying weather index can be directly modeled to estimate the fair price of derivative contracts. The IM relies on statistical and stochastic representations of the weather index to construct its probability distribution, from which the expected payoff is derived. Once the distribution of the weather index is identified and its parameters are estimated, simulated index values are generated to compute expected payoffs. Similar to the HBA method, the mean of these simulated payoffs represents the fair price of the contract.
@@ 4 Empirical Analysis
Each component of the ACI consists of a time series, either on a monthly basis, or for meteorological seasons (three-month periods ending in February, May, August, and November). These time series, starting in 1961, are derived from measurements collected at meteorological stations across the U.S. and Canada. The data sets are constructed separately for each country, in combination, and across 12 sub-regions. This study uses the U.S. data, where states are grouped into seven sub-regions based on their geographic locations: Alaska (ALA), Central East Atlantic (CEA), Central West Pacific (CWP), Midwest (MID), Southeast Atlantic (SEA), Southern Plains (SPL), and Southwest Pacific (SWP). The analyses use standardized seasonal values for each of the six ACI components (T90, T10, P, D, W, S) as well as the composite index (ACI). This seasonal data, obtained from the official Actuaries Climate IndexTM website [^1] ACI website: https://actuariesclimateindex.org/data/, cover a 64-year period from 1961 to 2024 for all seven sub-regions.
@@
Regarding weather-based indexes, the National Oceanic and Atmospheric Administration (NOAA) Climate database [^2] NOAA database: https://www.ncei.noaa.gov/access/monitoring/products/ provides monthly data on CDD, HDD, and PRE for all U.S. states, covering the period from 1895 to the present. To align with the ACI data, seasonal values for CDD, HDD, and PRE between 1961 and 2024 are calculated by summing the monthly values for each meteorological season. Sub-regional data is subsequently generated by averaging the seasonal values of all states within each region. However, due to the unavailability of CDD and HDD data for Alaska, the analyses are performed on the six remaining regions of the U.S.